Hi everyone, I need help on this word problem. I have been working on many word problems like the one below but this particular problem I can't seem to find the equation. I would really appreciate the help, thank you.

If the sum of a number and one is doubled the result is seven less than the number. Find the number.

Answer by mananth(16946)

let x be the number
2*(x+1)=x-7
2x+2=x-7
2x-x=-7-2
x=-9